OpenStudy (anonymous):

Can someone please help me in this question? Find (exactly) all solutions in [0, 2pie] for each equation: tan(t) = 1/tan(t)

OpenStudy (anonymous):

Use \[ \tan(t) =\frac{\sin(t)}{\cos(t)} \]

OpenStudy (anonymous):

Plug that into your equation, what do you get?

OpenStudy (anonymous):

Hint: There are exactly two solutions in the interval [0, 2Pi].

OpenStudy (anonymous):

Oh okay, so would I plug the 0,2Pi in t?

OpenStudy (anonymous):

? You don't

OpenStudy (anonymous):

hmm the answer is -.4995 but they don't seem to give me any reference like a graph or something :S

OpenStudy (anonymous):

lol the answer can impossibly be negative because the problem constraint is that t is in [0,2Pi]

OpenStudy (anonymous):

The answers are Pi/4 and 5Pi/4

OpenStudy (anonymous):

lol thats wierd, i'll look over it again, thanks for the help tho I appreciate it.

OpenStudy (anonymous):

You need to look for values t such that \[ \sin^2(t) = \cos^2(t) \]

OpenStudy (jamesj):

Alternatively, note that tan x = 1/tan x if and only if tan^2 x = 1, hence \[\tan x = \pm 1\] Look now at the graph of tan x on the interval [0, 2 pi] ...

OpenStudy (anonymous):

oh ic, thx